# Bulk Motions in the Black Hole Jet Sheath as a Candidate for the Comptonizing Corona

### v1

@@ Abstract
By means of two-dimensional general relativistic resistive magnetohydrodynamic simulations, we investigate the properties of the sheath separating the black hole jet from the surrounding medium. We find that the electromagnetic power flowing through the jet sheath is comparable to the overall accretion power, and is an important site of energy dissipation as revealed by the copious appearance of reconnection layers and plasmoid chains: $\sim 20\%$ of the radially flowing electromagnetic power is found to be dissipated between 2 and 10 $R_{\rm g}$. The plasma in these dissipation regions moves along a nearly paraboloidal surface with trans-relativistic bulk motions dominated by the radial component, whose dimensionless 4-velocity is $\sim 1.2\pm 0.5$. In the frame moving with the mean (radially-dependent) velocity, the distribution of stochastic bulk motions resembles a Maxwellian with an ‘effective bulk temperature’ of $\sim$100 keV. Scaling the global simulation to Cygnus X-1 parameters gives a rough estimate of the Thomson optical depth across the jet sheath $\sim 0.01-0.1$, and it may increase in future magnetohydrodynamic simulations with self-consistent radiative losses. These properties suggest that the dissipative jet sheath may be a viable ‘coronal’ region, capable of upscattering seed thermal X-ray photons ($\sim$1 keV) into a hard, nonthermal tail, as seen during the hard states of X-ray binaries and active galactic nuclei.
black hole physics – magnetic reconnection – relativistic processes – X-rays: binaries
@@ 1.1 Microphysics
Magnetic reconnection is incapable of sustaining a high electron temperature in the radiative corona (Beloborodov, 2017). It works in such a way that most of the plasma remains cool; however, the reconnection layer still efficiently Comptonizes photons by random bulk motions of the self-similar plasmoid chain created in the layer. This mechanism was found capable of producing the hard-state spectrum (Beloborodov, 2017) and then demonstrated by first-principle, radiative, kinetic plasma (particle-in-cell; PIC) simulations of magnetic reconnection (Sironi & Beloborodov 2020; Sridhar et al. 2021, 2023, hereafter, SB20, SSB21, SSB23, see also Werner et al. 2018; Mehlhaff et al. 2020). The bulk motions of plasmoids/magnetic islands are generated as a natural outcome of magnetic reconnection and can achieve trans-relativistic speeds (with an energy distribution resembling a Maxwellian having an ‘effective bulk temperature’ of $\sim 100$ keV). The stochastic bulk motions of this chain of plasmoids can Compton upscatter the soft X-ray photons to a hard, nonthermal tail. This was termed the cold-chain Comptonization process, as most of the electrons that reside inside the chain of Comptonizing plasmoids are cooled down to non-relativistic temperatures due to inverse Compton (IC) losses. A similar process can also operate in turbulent plasmas where the bulk motions of the eddies can Compton upscatter soft photons to a nonthermal distribution (Groselj et al. 2023; Nättilä 2024).
@@ 1.2 Macrophysics
X-ray spectral and temporal studies of black hole X-ray binaries have suggested that the location, size, and geometry of the corona could change depending on the phase of X-ray binary outburst (Kara et al., 2019; Sridhar et al., 2019, 2020; Wang et al., 2021; Cao et al., 2021; Méndez et al., 2022). Furthermore, the nonthermal X-ray emission during the hard states was typically seen contemporaneous with the presence of a compact ($\lesssim 10^{15}\,{\rm cm}$) radio jet (Fender & Kuulkers, 2001; Gallo et al., 2003; Fender et al., 2004; Gallo et al., 2018; Méndez et al., 2022). These observations have motivated models of coronae—albeit involving the less feasible thermal Comptonization process by hot electrons—involving the jet or its base (Markoff et al., 2005; Qiao & Liu, 2015; Lucchini et al., 2022). Recent results from the Imaging X-ray Polarimetry Explorer may also help constrain the corona geometry (IXPE; Weisskopf et al., 2016; Krawczynski et al., 2022; Marinucci et al., 2022; Ursini et al., 2022; Tagliacozzo et al., 2023).
@@
The corona was proposed to outflow with a mildly relativistic speed from the accretion disk in order to explain the hard spectral index and the weak reflection in hard-state black holes (Beloborodov, 1999; Malzac et al., 2001). In addition, Beloborodov (1998) showed that such outflows should generate polarization parallel to the disk normal. The picture of an outflowing corona has been used recently to explain the parallel polarization degree (PD) of $4.0\%\pm 0.2\%$ observed by IXPE in the 2–8 keV band from Cygnus X-1 (Poutanen et al., 2023). Dexter & Begelman (2024) further argued that the Comptonizing region might be a trans-relativistic outflow situated along the jet sheath. A similar picture has been proposed based on global radiative general relativistic magnetohydrodynamic (GRMHD) simulations (Moscibrodzka, 2023).
@@
Global GRMHD simulations showing potential regions of magnetic dissipation can provide insights into the geometry and orientation of a putative corona, as well as the spatial and temporal variations in the polarization properties of the Comptonized emission (Schnittman & Krolik, 2010; Beheshtipour et al., 2017). Within the tenuous corona, magnetic dissipation could occur in a collisionless regime (Goodman & Uzdensky, 2008). However, performing global PIC simulations of a realistic accreting system that would accurately capture the global dynamics and kinetic dissipation scales is computationally unfeasible (although two-dimensional general relativistic PIC simulations by Galishnikova et al. 2023 with simplifying assumptions on the accretion model, show that magnetic flux eruptions are a generic result of saturation of flux on the horizon). A simpler approach is based on GRMHD simulations with explicit resistivity (e.g., Qian et al. 2018; Vourellis et al. 2019; Ripperda et al. 2020). [^1] Although note that the magnetic reconnection rate in collisional plasma as described by MHD is 0.1$\sim$ 0.1$\times$ 0.1 × of the collisionless rate as found in PIC simulations Such simulations allow modeling the global dynamics of accretion disks while resolving reconnection in the largest current sheets with a converged dissipation rate. GRMHD simulations have revealed the formation of current sheets and plasmoid-unstable reconnection layers in two regions: (1) the sheath at the interface between the accretion disk and the Poynting-flux-dominated polar jet (Ripperda et al., 2020; Nathanail et al., 2020; Chashkina et al., 2021; Dihingia et al., 2022; Nathanail et al., 2022), and (2) the magnetospheric region between the black hole and the inner accretion disk (Ripperda et al., 2020, 2022) (See also a number of GRMHD and PIC simulations of the magnetospheric region without an accretion disk: Parfrey et al. 2019; Inda-Koide et al. 2019; Bransgrove et al. 2021; Hirotani et al. 2021; Crinquand et al. 2022; El Mellah et al. 2023).
@@
GRMHD simulations of accretion onto black holes including plasma-radiation interaction (Dexter et al., 2021; Liska et al., 2022) have shown the formation of a truncated inner thicker disk (and an outer thinner disk) with a strong poloidal field (akin to a MAD state) in case the disk is initially threaded with a dominant poloidal flux at large radii. Our simulation aims at modeling the properties of the inner regions of such a disk, including the current sheets, potential dissipation regions, and the fluid and electromagnetic field conditions found there. Unlike e.g., Liska et al. (2022), our simulations are non-radiative. While radiative cooling will strongly affect the conditions of the dense material accreting near the midplane, the properties of the jet and its sheath, being dominated by magnetic stresses, are conceivably similar between radiative and non-radiative models. So, simulations of accretion disks without radiative cooling could reasonably inform us about the magnetically dominated inner regions of the accretion flow during the low/hard state, in particular to understand the formation of current sheets and other dissipation regions found there. Therefore, the results from the simulation presented in this work can potentially capture some characteristics of the low/hard state dynamics and may serve to model the nonthermal X-rays seen during these phases.
The paper is organized as follows. In §2, we detail the numerical setup of the GRRMHD simulations. §3 contains the results, with §3.1 characterizing the jet sheath, §3.2 exploring the properties of the current sheets in the jet sheath, §3.3 focusing on the power flowing through the jet sheath, §3.4 investigating the properties of the bulk motions in this region, and §3.5 checking if the inferred optical depth in the jet sheath region that we identify as a corona satisfies the requirements to produce the observed X-ray spectra. Finally, we summarize our findings in §4.
@@ 2 GRRMHD simulation setup
The results presented in this work are obtained using two-dimensional GRRMHD simulations of prograde accretion onto a highly spinning black hole (Ripperda et al., 2020). The simulation was performed using the Black Hole Accretion Code (BHAC; Porth et al., 2017; Olivares et al., 2019) with implicit-explicit time-stepping scheme (Ripperda et al., 2019b) and adaptive mesh refinement (AMR) to accurately resolve dissipation sites on the resistive scale in the system. To accurately capture the regime of plasmoid formation and evolution (which is associated with magnetic dissipation), the Lundquist number $S=L_{\rm cs}v_{\rm A}/\eta$ should be above the threshold of $10^{4}$ (Loureiro et al., 2007; Bhattacharjee et al., 2009; Uzdensky et al., 2010). Taking a small, uniform plasma resistivity of $\eta=5\times 10^{-5}$ in the GRRMHD equations, since the Alfvén speed is $v_{\rm A}/c=\sqrt{\sigma/(\sigma+1)}\sim 1$ in the highly magnetized regions surrounding the jet ($\sigma\gtrsim 1$), the length scale of the resolved current sheets that would go plasmoid-unstable was found to be $L_{\rm cs}\gtrsim{\cal O}(1)\,R_{\rm g}$ (Ripperda et al., 2020), where $R_{\rm g}=GM_{\bullet}/c^{2}$ is the gravitational radius of a black hole with mass $M_{\bullet}$. The simulation results presented here thus reliably capture plasmoid formation in current sheets with length $\gtrsim R_{\rm g}$ near the black hole.
Figure 2: Panels [a,b,c] in the top row show the comoving plasma density, the dimensionless temperature ($p/\rho c^{2}$), and the bulk energy per unit rest mass energy ($\Gamma-1$), at time $Tc/R_{\rm g}=2870$. The black hole is centered at $[x,z]=[0,0]$, the red dashed curve is the ergosphere, and the red solid curve is the inner light surface. The overplotted curves are magnetic field lines. The time-evolution of these parameters (during the quasi-steady state, $Tc/R_{\rm g}>2000$) at a fiducial radius of $r=15\,R_{\rm g}$ is shown in the bottom row panels [d,e,f] as a function of the polar angle $\theta$ (the white dotted curves in the top row indicate the fiducial radius). The white left- and right-facing triangles in each of the bottom panels denote the time corresponding to the top panels. We only show the northern hemisphere, $\theta<\pi/2$.
@@
The base resolution of the simulation is $384\times 192$ (the two-dimensional nature of our simulation leaves the plasma dynamics along the $\phi$ coordinate invariant), and the simulation employs 5 levels of AMR, effectively corresponding to a resolution of 6144 radial cells and 3072 polar cells. The simulation is initiated from a torus in hydrodynamic equilibrium (with an equilibrium pressure distribution $p_{\rm eq}$; Fishbone & Moncrief, 1976) around a black hole with a dimensionless spin parameter $a_{\bullet}=0.9375$. The torus’s inner radius is at $r_{\rm in}=20\,R_{\rm g}$, and its density peaks at $r_{\rm max}=42\,R_{\rm g}$. The initiated torus is threaded by a weak poloidal magnetic field loop with a magnetic field strength set such that the plasma beta $\beta=2p/b^{2}=100$, where $p$ is the plasma pressure, and $b$ is the magnetic field strength in the frame comoving with the fluid (Bucciantini & Del Zanna, 2013). We also define a ‘cold magnetization’ $\sigma_{\rm c}\equiv b^{2}/(\rho c^{2})$ and a ‘hot magnetization’ $\sigma_{\rm h}\equiv b^{2}/(\rho hc^{2})$, where $\rho$ is the comoving plasma mass density and $h=1+4p/(\rho c^{2})$ is the specific enthalpy of an ideal gas with an adiabatic index of $\hat{\gamma}=4/3$. The atmosphere of the disk is permeated with diffuse plasma such that its Lorentz factor $\Gamma<20$, rest-mass density $\rho_{\rm atm}=\rho_{\rm min}r^{-3/2}$, and pressure $p_{\rm atm}=p_{\rm min}r^{-5/2}$, where $\rho_{\rm min}=10^{-4}$ and $p_{\rm min}=10^{-6}/3$ are the floor values of the density and pressure of the gas, respectively. The ceiling on the cold magnetization is set to be $\sigma_{\rm c,max}=100$. The magnetorotational instability (MRI; Velikhov, 1959; Chandrasekhar, 1960; Balbus & Hawley, 1991) is triggered by perturbing the equilibrium fluid pressure as $p=p_{\rm eq}(1+X_{\rm p})$ with a uniformly distributed random variable $X_{\rm p}\in[-0.02,0.02]$.

### v2

@@ Abstract
Using two-dimensional general relativistic resistive magnetohydrodynamic simulations, we investigate the properties of the sheath separating the black hole jet from the surrounding medium. We find that the electromagnetic power flowing through the jet sheath is comparable to the overall accretion power of the black hole. The sheath is an important site of energy dissipation as revealed by the copious appearance of reconnection layers and plasmoid chains. About 20% of the sheath power is dissipated between 2 and 10 gravitational radii. The plasma in the dissipative sheath moves along a nearly paraboloidal surface with trans-relativistic bulk motions dominated by the radial component, whose dimensionless 4-velocity is $\sim 1.2\pm 0.5$. In the frame moving with the mean (radially-dependent) velocity, the distribution of stochastic bulk motions resembles a Maxwellian with an ‘effective bulk temperature’ of $\sim$100 keV. Scaling the global simulation to Cygnus X-1 parameters gives a rough estimate of the Thomson optical depth across the jet sheath $\sim 0.01-0.1$, and it may increase in future magnetohydrodynamic simulations with self-consistent radiative losses. These properties suggest that the dissipative jet sheath may be a viable ‘coronal’ region, capable of upscattering seed soft photons into a hard, nonthermal tail, as seen during the hard states of X-ray binaries and active galactic nuclei.
Black hole physics (159) – high-energy astrophysics (739) – Magnetohydrodynamical simulations (1966) – Plasma astrophysics (1261) – X-ray binary stars (1811) – X-ray active galactic nuclei (2035)
@@ 1.1 Microphysics
Magnetic reconnection is incapable of sustaining a high electron temperature in the radiative corona (Beloborodov, 2017). It works in such a way that most of the plasma remains cool; however, the reconnection layer still efficiently Comptonizes photons by random bulk motions of the self-similar plasmoid chain created in the layer. This mechanism was found capable of producing the hard-state spectrum (Beloborodov, 2017) and then demonstrated by first-principle, radiative, kinetic plasma (particle-in-cell; PIC) simulations of magnetic reconnection (Sironi & Beloborodov 2020; Sridhar et al. 2021, 2023, hereafter, SB20, SSB21, SSB23, see also Werner et al. 2018; Mehlhaff et al. 2020). The bulk motions of plasmoids/magnetic islands are generated as a natural outcome of magnetic reconnection and can achieve trans-relativistic speeds (with an energy distribution resembling a Maxwellian having an ‘effective bulk temperature’ of $\sim 100$ keV). The stochastic bulk motions of this chain of plasmoids can Compton upscatter the soft X-ray photons to a hard, nonthermal tail. This was termed the cold-chain Comptonization process, as most of the electrons that reside inside the chain of Comptonizing plasmoids are cooled down to non-relativistic temperatures due to inverse Compton (IC) losses. A similar process can also operate in turbulent plasmas where the bulk motions of the eddies can Compton upscatter soft photons to a nonthermal distribution (Grošelj et al. 2024; Nättilä 2024).
@@ 1.2 Macrophysics
X-ray spectral and temporal studies of black hole X-ray binaries have suggested that the location, size, and geometry of the corona could change depending on the phase of X-ray binary outburst (Kara et al., 2019; Sridhar et al., 2019, 2020; Wang et al., 2021; Cao et al., 2021; Méndez et al., 2022). Furthermore, the nonthermal X-ray emission during the hard states was typically seen contemporaneously with the presence of a compact ($\lesssim 10^{15}\,{\rm cm}$) radio jet (Fender & Kuulkers, 2001; Gallo et al., 2003; Fender et al., 2004; Gallo et al., 2018; Méndez et al., 2022). These observations have motivated models of coronae—albeit involving the less feasible thermal Comptonization process by hot electrons—that involve the jet or its base (Markoff et al., 2005; Qiao & Liu, 2015; Lucchini et al., 2022). Recent results from the Imaging X-ray Polarimetry Explorer may also help constrain the coronal geometry (IXPE; Weisskopf et al., 2016; Krawczynski et al., 2022; Marinucci et al., 2022; Ursini et al., 2022; Tagliacozzo et al., 2023).
@@
The corona was proposed to outflow at a mildly relativistic speed from the accretion disk in order to explain the hard spectral index and the weak reflection in hard-state black holes (Beloborodov, 1999; Malzac et al., 2001). In addition, Beloborodov (1998) showed that such outflows should generate polarization parallel to the disk normal. The picture of an outflowing corona has been used recently to explain the parallel polarization degree (PD) of $4.0\%\pm 0.2\%$ observed by IXPE in the 2–8 keV band from Cygnus X-1 (Poutanen et al., 2023). Dexter & Begelman (2024) further argued that the Comptonizing region might be a trans-relativistic outflow situated along the jet sheath. A similar picture has been proposed based on global radiative general relativistic magnetohydrodynamic (GRMHD) simulations (Moscibrodzka, 2023).
@@
Global GRMHD simulations showing potential regions of magnetic dissipation can provide insights into the geometry and orientation of a putative corona, as well as the spatial and temporal variations in the polarization properties of the Comptonized emission (Schnittman & Krolik, 2010; Beheshtipour et al., 2017). Within the tenuous corona, magnetic dissipation could occur in a collisionless regime (Goodman & Uzdensky, 2008). However, performing global PIC simulations of a realistic accreting system that would accurately capture the global dynamics and kinetic dissipation scales is computationally unfeasible (although two-dimensional general relativistic PIC simulations by Galishnikova et al. 2023 with simplifying assumptions on the accretion model, show that magnetic flux eruptions are a generic result of saturation of flux on the horizon). A simpler approach is based on GRMHD simulations with explicit resistivity (e.g., Qian et al. 2018; Vourellis et al. 2019; Ripperda et al. 2020). [^1] Although note that the magnetic reconnection rate in collisional plasma as described by MHD is about a tenth of the collisionless rate as found in PIC simulations Such simulations allow modeling the global dynamics of accretion disks while resolving reconnection in the largest current sheets with a converged dissipation rate. GRMHD simulations have revealed the formation of current sheets and plasmoid-unstable reconnection layers in two regions: (1) the sheath at the interface between the accretion disk and the Poynting-flux-dominated polar jet (Ripperda et al., 2020; Nathanail et al., 2020; Chashkina et al., 2021; Dihingia et al., 2022; Nathanail et al., 2022), and (2) the magnetospheric region between the black hole and the inner accretion disk (Ripperda et al., 2020, 2022) (see also a number of GRMHD and PIC simulations of the magnetospheric region without an accretion disk: Parfrey et al. 2019; Inda-Koide et al. 2019; Bransgrove et al. 2021; Hirotani et al. 2021; Crinquand et al. 2022; El Mellah et al. 2023).
@@
GRMHD simulations of accretion onto black holes including plasma-radiation interaction (Dexter et al., 2021; Liska et al., 2022) have shown the formation of a truncated inner thicker disk (and an outer thinner disk) with a strong poloidal field (akin to a MAD state) if the disk is initially threaded with a dominant poloidal flux at large radii. Our simulation aims at modeling the properties of the inner regions of such a disk, including the current sheets, potential dissipation regions, and the fluid and electromagnetic field conditions found there. Unlike, e.g., Liska et al. (2022), our simulations are non-radiative. While radiative cooling will strongly affect the conditions of the dense material accreting near the midplane, the properties of the jet and its sheath, being dominated by magnetic stresses, are conceivably similar between radiative and non-radiative models. So, simulations of accretion disks without radiative cooling could reasonably inform us about the magnetically dominated inner regions of the accretion flow during the low/hard state, in particular to understand the formation of current sheets and other dissipation regions found there. Therefore, the results from the simulation presented in this work can potentially capture some characteristics of the low/hard state dynamics and may serve to model the nonthermal X-rays seen during these phases.
The paper is organized as follows. In §2, we detail the numerical setup of the GRRMHD simulations. §3 contains the results, with §3.1 characterizing the jet sheath, §3.2 exploring the properties of the current sheets in the jet sheath, §3.3 focusing on the power flowing through the jet sheath, §3.4 investigating the properties of the bulk motions in this region, and §3.5 checking whether the inferred optical depth in the jet sheath region that we identify as a corona satisfies the requirements to produce the observed X-ray spectra. Finally, we summarize our findings in §4.
@@ 2 GRRMHD simulation setup
The results presented in this work are obtained using two-dimensional GRRMHD simulations of prograde accretion onto a highly spinning black hole (Ripperda et al., 2020). The simulation was performed using the Black Hole Accretion Code (BHAC; Porth et al., 2017; Olivares et al., 2019) with an implicit-explicit time-stepping scheme (Ripperda et al., 2019b) and adaptive mesh refinement (AMR) to accurately resolve dissipation sites on the resistive scale in the system. To accurately capture the regime of plasmoid formation and evolution (which is associated with magnetic dissipation), the Lundquist number $S=L_{\rm cs}v_{\rm A}/\eta$ should be above the threshold of $10^{4}$ (Loureiro et al., 2007; Bhattacharjee et al., 2009; Uzdensky et al., 2010). Taking a small, uniform plasma resistivity of $\eta=5\times 10^{-5}$ in the GRRMHD equations, since the Alfvén speed is $v_{\rm A}/c=\sqrt{\sigma/(\sigma+1)}\sim 1$ in the highly magnetized regions surrounding the jet ($\sigma\gtrsim 1$), the length scale of the resolved current sheets that would go plasmoid-unstable was found to be $L_{\rm cs}\gtrsim{\cal O}(1)\,R_{\rm g}$ (Ripperda et al., 2020), where $R_{\rm g}=GM_{\bullet}/c^{2}$ is the gravitational radius of a black hole with mass $M_{\bullet}$. The simulation results presented here thus reliably capture plasmoid formation in current sheets with length $\gtrsim R_{\rm g}$ near the black hole.
Figure 2: Panels [a,b,c] in the top row show the comoving plasma density, the dimensionless temperature ($p/\rho c^{2}$), and the bulk energy per unit rest mass energy ($\Gamma-1$), at time $Tc/R_{\rm g}=2870$. The black hole is centered at $[x,z]=[0,0]$, the red solid curve is the ergosphere, and the red dashed curve is the inner light surface. The overplotted curves are magnetic field lines. The time-evolution of these parameters (during the quasi-steady state, $Tc/R_{\rm g}>2000$) at a fiducial radius of $r=15\,R_{\rm g}$ is shown in the bottom row panels [d,e,f] as a function of the polar angle $\theta$ (the white dotted curves in the top row indicate the fiducial radius). The white left- and right-facing triangles in each of the bottom panels denote the time corresponding to the top panels. We only show the northern hemisphere, $\theta<\pi/2$.
@@
The base resolution of the simulation is $384\times 192$ (the two-dimensional nature of our simulation leaves the plasma dynamics along the $\phi$ coordinate invariant), and the simulation employs five levels of AMR, effectively corresponding to a resolution of 6144 radial cells and 3072 polar cells. The simulation is initiated from a torus in hydrodynamic equilibrium (with an equilibrium pressure distribution $p_{\rm eq}$; Fishbone & Moncrief, 1976) around a black hole with a dimensionless spin parameter $a_{\bullet}=0.9375$. The torus’s inner radius is at $r_{\rm in}=20\,R_{\rm g}$, and its density peaks at $r_{\rm max}=42\,R_{\rm g}$. The initiated torus is threaded by a weak poloidal magnetic field loop with a magnetic field strength set such that the plasma beta $\beta=2p/b^{2}=100$, where $p$ is the plasma pressure, and $b$ is the magnetic field strength in the frame comoving with the fluid (Bucciantini & Del Zanna, 2013). We also define a ‘cold magnetization’ $\sigma_{\rm c}\equiv b^{2}/(\rho c^{2})$ and a ‘hot magnetization’ $\sigma_{\rm h}\equiv b^{2}/(\rho hc^{2})$, where $\rho$ is the comoving plasma mass density and $h=1+4p/(\rho c^{2})$ is the specific enthalpy of an ideal gas with an adiabatic index of $\hat{\gamma}=4/3$. The atmosphere of the disk is permeated with diffuse plasma such that its Lorentz factor $\Gamma<20$, rest-mass density $\rho_{\rm atm}=\rho_{\rm min}r^{-3/2}$, and pressure $p_{\rm atm}=p_{\rm min}r^{-5/2}$, where $\rho_{\rm min}=10^{-4}$ and $p_{\rm min}=10^{-6}/3$ are the floor values of the density and pressure of the gas, respectively. The ceiling on the cold magnetization is set to be $\sigma_{\rm c,max}=100$. The magnetorotational instability (Velikhov, 1959; Chandrasekhar, 1960; Balbus & Hawley, 1991) is triggered by perturbing the equilibrium fluid pressure as $p=p_{\rm eq}(1+X_{\rm p})$ with a uniformly distributed random variable $X_{\rm p}\in[-0.02,0.02]$.